OpenStudy (keke):

simplify; √(18)+√(108)+√(50)

OpenStudy (angela210793):

3sqr2+6sqr3+5sqr2=8sqr2+6sqr3

OpenStudy (angela210793):

do u understand it?

OpenStudy (anonymous):

\[\sqrt{18}+\sqrt{108}+\sqrt{50} \]?

OpenStudy (anonymous):

\[3\sqrt{2}+6\sqrt{3}+5\sqrt{2}=8\sqrt{2}+6\sqrt{3}\]